OpenStudy (anonymous):

How do you simplify: [9-x^(-2)] / [3+x^(-1)]

OpenStudy (anonymous):

well i can tell you that x^2-2= 1/x^2 in general x^-n=1/x^n hope that helps

OpenStudy (anonymous):

(9+x)/(3-x²)?

OpenStudy (anonymous):

(3*x-1) --------- x

OpenStudy (anonymous):